OpenStudy (anonymous):

Solve the equation for 0

OpenStudy (nubeer):

first divide both sides by 6 2) take square root on both sides.

sam (.sam.):

What do you get @ChongVinny

OpenStudy (anonymous):

@.Sam. 65.55. why ?

OpenStudy (anonymous):

@nubeer thanks alot ! :)

sam (.sam.):

You still have other values

sam (.sam.):

\[\sin^{2}(x)=\frac{5}{6} \\ \\ \sin(x)=\pm \sqrt{\frac{5}{6}} \\ \\ x= \pm \sin^{-1}(\sqrt{\frac{5}{6}})\]

OpenStudy (anonymous):

so another one is negative ?

sam (.sam.):



sam (.sam.):

We get x=65.9 Then we got plus and minus for sine, so 180-65.9=114.1 180+65.9=245.9 360-65.9=294.1 But 0<x<180 So x=65.9,114.1 degrees

OpenStudy (anonymous):

yayy ! i understand already. thanks ! :))

OpenStudy (dan815):